OpenStudy (anonymous):

use implicit differentiation to find dy/dx:

OpenStudy (anonymous):

\[\cos(r)+\cot(\theta)=e ^{r \theta}\]

OpenStudy (anonymous):

Please help! I'm so lost in my calc class and I don't start with my tutor until tomorrow.. :(

OpenStudy (anonymous):

You want to find dy/dx or dr/dtheta or dtheta/dr?

OpenStudy (anonymous):

dr/dtheta

OpenStudy (anonymous):

Okay: So firstly you want to differentiation both sides: \[\frac{d}{d \theta}(\cos(r)+\cot(\theta))=\frac{d}{d \theta}e^{r \theta}\] So for the first term: \[\frac{d}{d \theta}\cos(r)=-\sin(r) * \frac{d r }{d \theta}\] For the second term there is no "r" so its a "normal" derivative: \[\frac{d}{d \theta}\cot(\theta)=-\csc^2(\theta)\] Now the hardest term is the right side. When you use the chain rule you differentiate the outside, and the inside (via the chain rule). So you would have: \[\frac{d}{d \theta}e^{r \theta}=e^{r \theta} * \frac{d}{d \theta}(r \theta)=e^{r \theta}(\frac{dr}{d \theta} \theta+r (1))\] implies: \[-\sin(r) \frac{dr}{d \theta}-\csc^2(\theta)=e^{r \theta}(\frac{dr}{d \theta} \theta+r)\] Now you need simply solve for dr/d theta

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

No problem!